8.A computer assembling company receives 24% of parts from supplier X, 36% of parts from supplier Y, and the remaining 40% of parts from supplier Z. Five percent of parts supplied by X, ten percent of parts supplied by Y, and six percent of parts supplied by Z are defective. If an assembled computer has a defective part in it, what is the probability that this part was received from supplier Z?
Answer by robertb(5830)

P(Z|def) = P(def AND Z)/P(def)
=(P(def|Z)*P(Z))/(P(def|X)*P(X) + P(def|Y)*P(Y) + P(def|Z)*P(Z))
= 0.06*0.40/(0.05*0.24+0.10*0.36+0.06*0.40) = 0.024/0.072 = 1/3